McCullough Hospital uses a job-order costing system to assign costs to its patients. Its direct materials include a variety of items such as pharmaceutical drugs, heart valves, artificial hips, and pacemakers. Its direct labor costs (e.g., surgeons, anesthesiologists, radiologists, and nurses) associated with specific surgical procedures and tests are traced to individual patients. All other costs, such as depreciation of medical equipment, insurance, utilities, incidental medical supplies, and the labor costs associated with around-the-clock monitoring of patients are treated as overhead costs.

Historically, McCullough has used one predetermined overhead rate based on the number of patient-days (each night that a patient spends in the hospital counts as one patient-day) to allocate overhead costs to patients. Recently a member of the hospital’s accounting staff has suggested using two predetermined overhead rates (allocated based on the number of patient-days) to improve the accuracy of the costs allocated to patients. The first overhead rate would include all overhead costs within the Intensive Care Unit (ICU) and the second overhead rate would include all Other overhead costs. Information pertaining to the hospital’s estimated number of patient-days, its estimated overhead costs, and two of its patients—Patient A and Patient B—is provided below:

ICU Other Total
Estimated number of patient-days 2,600 5,200 7,800
Estimated fixed overhead cost $ 4,651,400 $ 4,248,400 $ 8,899,800
Estimated variable overhead cost per patient-day $ 239 $ 110
Patient A Patient B
Direct materials $ 5,100 $ 6,800
Direct labor $ 26,500 $ 37,200
Total number of patient-days (including ICU) 15 20
Number of patient-days spent in ICU 0 8

Required:

1. Assuming McCullough uses only one predetermined overhead rate, calculate:

a. The predetermined overhead rate.

b. The total cost, including direct materials, direct labor and applied overhead, assigned to Patient A and Patient B.

2. Assuming McCullough calculates two overhead rates as recommended by the staff accountant, calculate:

a. The ICU and Other overhead rates.

b. The total cost, including direct materials, direct labor and applied overhead, assigned to Patient A and Patient B.
Answer #1
1a
Estimated fixed overhead cost 8899800
Estimated variable overhead cost 1193400 =(2600*239)+(5200*110)
Estimated total overhead cost 10093200
Divide by Total Estimated number of patient-days 7800
The predetermined overhead rate 1294 per patient day
1b
Patient A
Direct materials 5100
Direct labor 26500
Applied overhead 19410 =15*1294
Total cost for Patient A 51010
Patient B
Direct materials 6800
Direct labor 37200
Applied overhead 25880 =20*1294
Total cost for Patient B 69880
2a
Estimated fixed overhead cost 4651400
Estimated variable overhead cost 621400 =2600*239
Estimated total overhead cost 5272800
Divide by Estimated number of patient-days 2600
Predetermined ICU overhead rate 2028 per patient day
Estimated fixed overhead cost 4248400
Estimated variable overhead cost 572000 =5200*110
Estimated total overhead cost 4820400
Divide by Estimated number of patient-days 5200
Predetermined Other overhead rate 927 per patient day
2b
Patient A
Direct materials 5100
Direct labor 26500
Applied overhead 13905 =15*927
Total cost for Patient A 45505
Patient B
Direct materials 6800
Direct labor 37200
Applied overhead 27348 =(8*2028)+(12*927)
Total cost for Patient B 71348